OpenStudy (lilkg77):

The first quartile of a data set is 41, and the third quartile is 57. Which of these values in the data set is an outlier? A. 83 B. 39 C. 19 D. 73

OpenStudy (lilkg77):

@shaik0124 @reemii

OpenStudy (lilkg77):

@i0iz0gangster

OpenStudy (reemii):

The definition of outlier has something to do with IQR ("InterQuartileRange"). Do you remember?

OpenStudy (shaik0124):

i don't know about outlier but ("InterQuartileRange") is q3-q1

OpenStudy (reemii):

q3-q1 = 57 - 41 = 16 = IQR An outlier is a number smaller than 41 - 1.5 IQR or greater than 57+1.5IQR. that is, smaller than 41 - 24 or greater than 57 + 24